OpenStudy (anonymous):

Solve for X and Y -3y = 2x - 9

OpenStudy (anonymous):

So this slope-intercept form is y = mx + b y=2/3x−3 So y = -3 and use the slope 2/3 to determine what x is Meaning graph the above slope-intercept form to get your answer :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Yuppers :P